if vector a+ vector b is perpendicular to vector b and vector + 2b  is perpendicular to vector a then

Grade:11

4 Answers

Zameer Ansari
6 years ago
The question is not clear..If any one could right the question properly it could be solved....try writing in the form as take a` to be vector a
Akhileshsingh
6 years ago
The dot product of vector a+ vector with vector b will equal to zero similarly dot product of vector 2 a with vector b is equal to zero
Arin Som
6 years ago
(a+b).b = 0 (because they are perpendicular)a.b + |b|² = 0. --> (1)(2b+a).a=02b.a + |a|² =0 a.b = -|a|²/2. (a.b=b.a) Put this eqn in eqn (1) and you`ll get your answer as |a|=√2|b|Hope this helps!
Priyanka
5 years ago
Given,
A+B=R→1
2A +B is  perpendicular to A
So
(2A+B)×A=0
2A²+AB=0
A=-2B→2
Substitute equation 2 in 1
A+B=R
-2B+B=R
B=R